### v1

# Multiplicity of powers of path ideals of a line graph

Abstract.
Let $S=K[x_{1},\ldots,x_{n}]$ and let $I$ be the $t$-path ideal of the line graph $L_{n}$ with $n$-vertices. It is shown that the set of associated prime ideals of $I^{s}$ is equal to the set of minimal prime ideals of $I$ for all $s\geq 1$, and we provide an explicit description of these prime ideals. Additionally, as the main contribution of this paper, we derive an explicit formula for the multiplicity of $S/I^{s}$ for all $s\geq 1$, revealing that it is a polynomial in $s$ from the beginning.
Key words and phrases:
path ideals, line graph, multiplicity, normally torsion-free
2010 Mathematics Subject Classification:
Primary 13C70, 13H10 Secondary 05E40
1. Introduction
The edge ideal was introduced and studied by Villarreal in [18]. Let $G$ be a finite simple graph with vertex set $V=\{x_{1},\cdots,x_{n}\}$ and edge set $E$. The ideal generated by all quadratic monomials $x_{i}x_{j}$ such that $\{x_{i},x_{j}\}\in E$ is called the edge ideal of $G$, denoted by $I(G)$. Since then, the study of edge ideals of simple graphs, particularly the properties of their powers, has become a vibrant and active topic in combinatorial commutative algebra. For more detailed information, refer to the exposition paper [2] and the references therein.
In 1999, Conca and De Negri introduced the concept of a $t$-path ideal in their work [7], which serves as a generalization of the definition of an edge ideal. For an integer $2\leq t\leq n$, a $t$-path of the graph $G$ is a sequence of $(t-1)$ distinct edges $\{x_{i_{1}},x_{i_{2}}\},\{x_{i_{2}},x_{i_{3}}\},\cdots,$$\{x_{i_{t-1}},x_{i_{t}}\}$ of $G$ such that the vertices $x_{i_{1}},x_{i_{2}},\ldots,x_{i_{t}}$ are pairwise distinct. Such a path is also denoted by $\{x_{i_{1}},\ldots,x_{i_{t}}\}$ for short. The $t$-path ideal $I_{t}(G)$ associated to $G$ is defined as the squarefree monomial ideal
$I_{t}(G)=(x_{i_{1}}\cdots x_{i_{t}}\colon\;\{x_{i_{1}},\ldots,x_{i_{t}}\}\textrm{~{}~{}is~{}~{} a}~{}~{}t\textrm{-path~{}~{} of}~{}~{}G)$
in the polynomial ring $K[x_{1},\ldots,x_{n}]$. In the recent years, some algebraic properties of path ideals have been investigated extensively, see for instance [1, 4, 5, 13, 15, 16].
[…]
In this paper, we will study algebraic properties of powers of $I_{t}(L_{n})$ from two different angles: associated prime ideals of $I_{t}(L_{n})^{s}$ and the multiplicities of $I_{t}(L_{n})^{s}$ for all $s\geq 1$.
1.1. Associated prime ideals
Let $R$ be a Noetherian ring and $M$ an finitely generated non-zero $R$-module. A prime ideal $\operatorname{\mathfrak{p}}$ of $R$ is called an associated prime ideal of $M$ if there exists $a\in M$ such that $\operatorname{\mathfrak{p}}=(0:_{R}a)$. The set of all associated prime ideal of $M$ is denoted by $\mathrm{Ass}_{R}(M)$. It is known that $\mathrm{Ass}_{R}(M)$ is a non-empty and finite set. If $I$ is an ideal of $R$, we denote $\mathrm{Ass}_{R}(R/I)$ by $\mathrm{ass}(I)$. In 1979, Brodmann [6] proved the remarkable result that there exists an integer $s_{0}$ such that $\mathrm{ass}(I^{s})=\mathrm{ass}(I^{s_{0}})$ for all $s\geq s_{0}$. This result serves as a good example of the concept in commutative algebra that ideals exhibit favorable behavior asymptotically. (Another example illustrating this idea is provided later in the paper regarding the multiplicity of an ideal.) Naturally, Brodmann’s theorem also inspires several questions, such as determining the value of $s_{0}$ and determine the prime ideals in $\mathrm{ass}^{\infty}(I)$? Here, $\mathrm{ass}^{\infty}(I):=\mathrm{ass}(I^{s_{0}})$. We refer to [8] and the references therein for a comprehensive introduction to this topic. If $s_{0}=1$, then $I$ is said to be normally torsion-free. More formally, we give the following definition following [10].
Definition 1.1.
An ideal $I$ of a Noetherian ring $R$ is said to be normally torsion-free if $\mathrm{ass}(I^{s})=\mathrm{ass}(I)$ for all $s\geq 1$.
We will show that both $I_{t}(L_{n})$ and $I_{t}(L_{n})^{\vee}$ are normally torsion-free and determine the prime ideals in $\mathrm{ass}^{\infty}(I_{t}(L_{n}))$.
1.2. Multiplicity
Let $S=K[x_{1},\ldots,x_{n}]$ be a polynomial ring that standard graded over a field $K$ and let $M$ be a finitely generated graded $S$-module. The Hilbert function $H_{M}$ is defined as the function
$H_{M}(k):=\dim_{\mathbb{K}}M_{k},$
where $M_{k}$ is the degree $k$ component of $M$. The Hilbert series $\operatorname{HS}(M,z)$ is defined as
$\operatorname{HS}(M,z):=\sum_{i\in\mathbb{Z}}\dim_{\mathbb{K}}M_{i}z^{i}.$
If we assume that $\dim M=d$, it is a celebrated result of David Hilbert that $\operatorname{HS}(M,z)$ is a rational function of the following form
$\operatorname{HS}(M,z)=\frac{\mathbf{Q}(M,z)}{(1-z)^{d}}.$
Here, $\mathbf{Q}(M,z)\in\mathbb{Q}[z^{-1},z]$ is a Laurent polynomial such that $\mathbf{Q}(M,1)\neq 0$. As a consequence of this, the Hilbert function $H_{M}$ is of polynomial type, meaning that there exists a polynomial $p_{M}\in\mathbb{Q}[x]$ such that $H_{M}(k)=p_{M}(k)$ for all $k\gg 0$. The polynomial $p_{M}$ is referred to as the Hilbert polynomial of $M$.
Definition 1.2.
Let $M$ be a finitely generated graded $S$-module of dimension $d+1$. The Hilbert polynomial $P_{M}(z)$ of $M$ can be written as
$P_{M}(z)=\sum_{i=0}^{d}(-1)^{i}e_{i}(M)\binom{z+d-i}{d-i}.$
The integer coefficients $e_{i}(M)$ are called the Hilbert coefficients of $M$. The first Hilbert coefficient $e_{0}(M)$ holds a special significance, it is also known as the multiplicity of $M$ and is often denoted by $\operatorname{mult}(M)$.
Let $I$ be a graded ideal of $S$ such that $\dim M/I^{s}M=d$. It was shown in [12, Theorem 1.1] that $e_{i}(M/I^{s}M)$ is of polynomial type in $s$ of degree $\leq n-d+i$ for $i=1,2,\ldots,d$. Recall that a function $f:\mathbb{N}\rightarrow\mathbb{Q}$ is of polynomial type of degree $d$ if there exists a polynomial $p\in\mathbb{Q}[z]$ of degree $d$ such that $f(n)=p(n)$ for all $n\gg 0$. To the best of our knowledge, there are currently no known non-trivial examples of graded ideals for which the multiplicity of their powers has been explicitly computed. In this paper, we will provide an explicit formula for $\operatorname{mult}(S/I_{t}(L_{n})^{s})$.
1.3. Structure of this paper
Let $I$ denote the $t$-path ideal of the line graph $L_{n}$. This paper is organized as follows. In Section 2 we demonstrate that both $I$ and its Alexander dual $I^{\vee}$ are normally torsion-free and provide a complete characterization of all the ideals in $\mathrm{ass}^{\infty}(I)$. Additionally, we establish that $S/I$ is sequentially Cohen-Macaulay and determine its projective dimension. In the final section, which constitutes the main contribution of this paper, we present an explicit formula for the multiplicity of $S/I^{s}$ for all $s\geq 1$, revealing that it is a polynomial in $s$ from the beginning.
1.4. Conventions
In the paper, we always assume $n,t$ are integers with $n\geq t\geq 1$. The polynomial ring $K[x_{1},\ldots,x_{n}]$ will be denoted by $S$. We denote $L_{n}$ as the line graph of length $n-1$ and $I_{t}(L_{n})$ as the $t$-path ideal of $L_{n}$. Let $G(I_{t}(L_{n}))$ denote the minimal monomial generating set of $I_{t}(L_{n})$. Specially, $G(I_{t}(L_{n}))=\{u_{1},\ldots,u_{n-t+1}\}$, where $u_{i}=x_{i}x_{i+1}\cdots x_{i+t-1}$ for $i=1,\ldots,n-t+1$. For a positive integer $a$, the set $\{1,\ldots a\}$ is denoted by $[a]$. Finally, for a finite set $A$, $|A|$ refers to the number of $A$.
We refer to the book [10] for the notation that is not explained in this paper.
2. Associated primes of $I_{t}(L_{n})^{s}$
Let $I$ denote the $t$-path ideal $I_{t}(L_{n})$ of the line graph $L_{n}$. In this section, we show that both $I$ and its Alexander dual $I^{\vee}$ are normally torsion-free and provide an explicit description of minimal prime ideals of $I$. Furthermore, We show that $S/I$ is sequentially Cohen-Macaulay and determine its projective dimension.
Recall that a simplicial complex on $[n]:=\{1,\ldots,n\}$ is a collection $\Delta$ of subsets of $[n]$ such that if $F\in\Delta$ and $G\subseteq F$, then $G\in\Delta$. Any $F\in\Delta$ is a face of $\Delta$. A maximal face (with respect to inclusion) is a facet of $\Delta$. The set of all facets of $\Delta$ is denoted by $\mathcal{F}(\Delta)$. A simplicial complex $\Delta$ is denoted by $\langle F_{1},F_{2},\ldots,F_{r}\rangle$ if $F_{1},F_{2},\ldots,F_{r}$ are all facets of $\Delta$.
For a subset $F$ of $[n]$, let $x_{F}$ denote the monomial $\prod_{i\in F}x_{i}$. Then the facet ideal of $\Delta$ is defined to be the monomial ideal
$I(\Delta)=(x_{F}\colon\;F\mbox{ is a facet of }\Delta),$
and the Stanley Reisner $I_{\Delta}$ ideal of $\Delta$ is defined as
$I_{\Delta}=(x_{F}\colon\;F\subseteq[n]\mbox{ and }F\notin\Delta).$
We may look upon $I_{t}(L_{n})$ as the facet ideal of some simplicial complex. For this, we put
$F_{i}:=\{i,\ldots,i+t-1\}\subseteq[n]\mbox{ for }i=1,\ldots,n-t+1$
and let $\Delta_{n,t}$ denote the simplicial complex $\langle F_{1},\ldots,F_{n-t+1}\rangle$. It is immediate to see that the path ideal $I_{t}(L_{n})$ is the facet ideal $I(\Delta_{n,t})$ of $\Delta_{n,t}.$ The following lemma reveals an interesting property of $\Delta_{n,t}$, which we will not use in our subsequent discussion.
Recall a simplicial complex $\Delta$ is called a flag complex if its Stanley-Reisner ideals is generated by monomials of degree 2, or equivalently, all minimal non-faces of $\Delta$ have exactly two elements. Here, a non-face of $\Delta$ refers to a subset of $[n]$ which does not belong to $\Delta$.
Lemma 2.1.
For all $2\leq t\leq n$, the Stanley-Reisner ideal $I_{\Delta_{n,t}}$ is generated by all quadratic monomials $x_{i}x_{j}$ with $j-i\geq t.$ In particular, $\Delta_{n,t}$ is a flag complex.
Proof.
Let $I$ denote the ideal $(x_{i}x_{j}\colon\;j-i\geq t).$ It is clear that if $j-i\geq t$, then $\{i,j\}$ is a non-face of $\Delta_{n,t}$. Thus, $I\subseteq I_{\Delta_{n,t}}$.
Conversely, let $\{i_{1},\ldots,i_{k}\}$ be a non-face of $\Delta_{n,t}$ with $i_{1}<\cdots<i_{k}$. Then if $i_{k}-i_{1}<t$, then $\{i_{1},\ldots,i_{k}\}\subseteq F_{i_{1}}$, a contradiction. This implies $i_{k}-i_{1}\geq t$ and so $x_{F}\in I$. This proves $I\supseteq I_{\Delta_{n,t}}.$
Recall that a facet $F$ of a simplicial complex $\Delta$ is called a leaf if either $F$ is the only facet of $\Delta$, or there exists a facet $G\neq F$ such that $H\cap F\subseteq G\cap F$ for any facet $H\neq F$. A leaf $F$ of $\Delta$ is said to be good if the sets $F\cap H$, where $H\in\mathcal{F}(\Delta)$ is linearly ordered. A simplicial complex $\Delta$ is a simplicial forest if every nonempty subcollection of $\Delta$ has a leaf. A connected simplicial forest is called a simplicial tree. Here, a subcollection of $\Delta$ refers to a simplicial complex whose set of facets is a subset of $\mathcal{F}(\Delta).$ The following observation forms the foundation of all the results in this section.
Lemma 2.2.
For all $0<t\leq n$, $\Delta_{n,t}$ is a simplicial tree.
Proof.
It is clear that $\Delta_{n,t}$ is connected. Let $\Delta^{\prime}=\langle F_{k_{1}},\ldots F_{k_{r}}\rangle$ be a subcollection of $\Delta$. Note that $F_{k_{i}}=\{k_{i},k_{i}+1,\ldots,k_{i}+t-1\}$ for $i=1,\ldots,r$. We may assume that $r\geq 2$ and that $11\leq k_{1}<\cdots<k_{r}\leq n-t+$ 1. It can be easily verified that
$F_{k_{1}}\cap F_{k_{i}}=\left\{\begin{array}[]{ll}\{k_{i},k_{i}+1,\ldots,k_{1}+t-1\},&\hbox{if $k_{i}\leq k_{1}+t-1$;}\\ \emptyset,&\hbox{otherwise}\end{array}\right.$
for $i=2,\ldots,r$. From this it follows that
$F_{k_{1}}\cap F_{k_{2}}\supset F_{k_{1}}\cap F_{k_{3}}\supset\cdots\supset F_{k_{1}}\cap F_{k_{r}}.$
Therefore $F_{k_{1}}$ is a (good) leaf of $\Delta^{\prime}$. Hence, $\Delta_{n,t}$ is a simplicial tree.
We now present the first main result of this section.
Theorem 2.3.
For all $1\leq t\leq n$, both $I_{t}(L_{n})$ and $I_{t}(L_{n})^{\vee}$ are normally torsion-free.
Proof.
It is shown in [14, Theorem 2.2 and Theorem 2.5] that if a simplicial complex $\Delta$ has no special odd cycles, then the facet ideal $I(\Delta)$ and its Alexander dual $I(\Delta^{*})=I(\Delta)^{\vee}$ are normally torsion-free. On the other hand, by [14, Theorem 3.2], if $\Delta$ is a simplicial forest, then $\Delta$ has no special cycles. Now, the result follows by Lemma 2.2.
By applying [10, Theorem 1.4.6], we immediately obtain the following corollary.
Corollary 2.4.
Let $1\leq t\leq n$. Then $I_{t}(L_{n})$ is normal, this is, $\overline{I_{t}(L_{n})^{s}}=I_{t}(L_{n})^{s}$ for all $s\geq 1$.
Corollary 2.5.
Let $\leq t\leq n$. Then the Rees algebra $\mathcal{R}(I_{t}(L_{n}))$ is a normal Cohen-Macauly domain.
Proof.
By [14, Corollary 1.6].
In the following, we will determine all the minimal prime ideals of $I_{t}(L_{n})$, which will be also useful in the next section. To do this, we introduce some other notation. If $F\subseteq[n]$, we denote by $\operatorname{\mathfrak{p}}_{F}$ the prime ideal generated by $x_{i}$ with $i\in F$.
Notation 2.6.
Let $C_{n,t}$ denote the set of subsets $\{i_{1},\ldots,i_{r}\}$ of $[n]$ with $1\leq i_{1}<\cdots<i_{r}\leq n$ and the following conditions are satisfied:
  1. (1)
    $1\leq i_{1}\leq t$;
  2. (2)
    $i_{2}>t$;
  3. (3)
    $1\leq i_{j+1}-i_{j}\leq t$ for all $j=1,\ldots,r-1$;
  4. (4)
    $i_{j+2}-i_{j}>t$ for all $j=1,\ldots,r-2$;
  5. (5)
    $i_{r-1}<n-t+1$;
  6. (6)
    $n-t+1\leq i_{r}\leq n$.
Recall a vertex cover of the simplicial complex $\Delta$ is a subset $C\subset[n]$ such that $C\cap F\neq\emptyset$ for all $F\in{\mathcal{F}}(\Delta)$. The vertex cover $C$ is called a minimal vertex cover, if no proper subset of $C$ is a vertex cover. It is well-known that $\operatorname{\mathfrak{p}}$ is a minimal prime ideal of $I(\Delta)$ if and only if $\operatorname{\mathfrak{p}}=\operatorname{\mathfrak{p}}_{C}$ for some minimal vertex cover of $\Delta$.
Theorem 2.7.
For $1\leq t\leq n$, the set of minimal prime ideals of $I_{t}(L_{n})$ is given by
$\{\operatorname{\mathfrak{p}}_{F}\colon\;F\in C_{n,t}\}.$
Proof.
It is equivalent to showing that $C_{n,t}$ consists of all minimal vertex cover of $\Delta_{n,t}$. Let $F\in C_{n,t}$. We first show that $F$ is a vertex cover of $\Delta_{n,t}$. Given a facet $F_{i}$ of $\Delta_{n,t}$, where $11\leq i\leq n-t+$ 1. If $i\in F$, then $i\in F\cap F_{i}$ and so $F\cap F_{i}\neq\emptyset$. Suppose that $i\notin F$. If $i<i_{1}$ then $i_{1}\in F_{i}\cap F$ and so $F_{i}\cap F\neq\emptyset$. If $i>i_{1}$, then, since $i_{r}\geq n-t+1$, there exists $11\leq j\leq r-$ 1 such that $i_{j}<i<i_{j+1}$. Since $i_{j+1}-i<i_{j+1}-i_{j}\leq t$, it follows that $i_{j+1}\in F_{i}$ and thus $F\cap F_{i}\neq\emptyset$. This proves $F$ is a vertex cover of $\Delta_{n,t}$. Next, we show that $F$ is minimal, that is, $F\setminus\{i_{j}\}$ is not a vertex cover of $\Delta_{n,t}$ for $j=1,\ldots,r$. But, this is clear by noticing that $F_{1}\cap(F\setminus\{i_{1}\})=\emptyset$, $F_{i_{j-1}+1}\cap(F\setminus\{i_{j}\})=\emptyset$ for $j=2,\ldots,r-1$ and $F_{n-t+1}\cap(F\setminus\{i_{r}\})=\emptyset$. Therefore, $F$ is a minimal vertex cover of $\Delta_{n,t}$ indeed.
Conversely, let $F$ be a minimal vertex cover of $\Delta_{n,t}$. We may write $F=\{i_{1},\ldots,i_{r}\}$ with $i_{1}<\ldots<i_{r}$. Since $F_{1}\cap F\neq\emptyset$, it follows that $i_{1}\leq t$ and so (1) holds. Similarly, since $F_{n-t+1}\cap F\neq\emptyset$, we have $i_{r}\geq n-t+1$ and this proves (6). If $i_{2}\leq t$, then, note that $i_{1}\in F_{i}$ implies $i_{2}\in F_{i}$ for all $i=1,\ldots,n-t+1$, it follows that $F\setminus\{i_{1}\}$ is also a vertex cover of $\Delta_{n,t}$, a contradiction. This establishes (2). Similarly, (5) has to hold. If $i_{j+1}-i_{j}>t$ for some $j\in\{1,\ldots,r-1\}$, then $F_{i_{j}+1}\cap F=\emptyset$. This is impossible and so (3) should hold. It remains to prove (4). Suppose on the contrary that $i_{j+2}-i_{j}\leq t$ for some $21\leq j\leq r-$ 2. Then, if $i_{j+1}\in F_{i}$ for some $i$, we have $i_{j+1}\in\{i,i+1,\ldots,i+t-1\}$. It follows that either $i_{j+2}\leq i+t-1$ (i.e., $i_{j+2}\in F_{i}$) or $i_{j}\geq i$ (i.e., $i_{j}\in F_{i}$), for otherwise $i_{i+2}-i_{j}>t$. Thus, $F\setminus\{i_{j+1}\}$ is also a vertex cover of $\Delta_{n,t}$, which is a contradiction. This completes the proof.
Corollary 2.8.
Let $2\leq t\leq n$. Then $\mathrm{ass}(I_{t}(L_{n})^{s})=\{\operatorname{\mathfrak{p}}_{F}\colon\;F\in C_{n,t}\}$ for all $s\geq 1$.
Corollary 2.9.
The Alexander dual $I_{t}(L_{n})^{\vee}$ of $I_{t}(L_{n})$ has the minimal generating set $G=\{X_{F}\colon\;F\in C_{n,t}\}$. Assume $n=(t+1)p+q$ with $0\leq q\leq t$. Then the largest degree of minimal generators of $I_{t}(L_{n})^{\vee}$ is given by
$\operatorname{Deg}(I_{t}(L_{n})^{\vee})=\left\{\begin{array}[]{ll}2p,&\hbox{$q<t$;}\\ 2p+1,&\hbox{$q=t$.}\end{array}\right.$
Proof.
Denote $\operatorname{Deg}(I_{t}(L_{n})^{\vee})$ as $D$. Consider the following infinite sequence
$1,t+1,t+2,2t+2,2t+3,\cdots,kt+k,kt+k+1,\cdots.$
Let $a_{m}$ denote the $m$-th number of this sequence. Namely, $a_{1}=1,a_{2}=t+1,a_{3}=t+2$ and so on. Thus, $a_{2k}=kt+k$ and $a_{2k+1}=kt+k+1$ for $k\geq 1$. Given $F=\{i_{1},\ldots,i_{r}\}\in C_{n,t}$ with $1\leq i_{1}<\cdots<i_{r}\leq n$, we claim that
$a_{m}\leq i_{m}\mbox{ for all }m=1,\ldots,r.$
In fact, $a_{1}=1\leq i_{1}$ and $a_{2}=t+1\leq i_{2}$. Assume that $12\leq m\leq r-$ 1 and $a_{j}\leq i_{j}$ for $j=1,\ldots,m$. Then, if $m=2k$ for some $k\geq 1$, then $a_{m+1}=a_{m}+1$ and so $i_{m+1}\geq i_{m}+1\geq a_{m}+1=a_{m+1}$; if $m=2k+1$, then $a_{m+1}-a_{m-1}=t+1$. Since $i_{m+1}-i_{m-1}\geq t+1$, it follows that $a_{m+1}\leq i_{m+1}$. This proves the claim.
We next consider the following cases:
Case 1: If $n=(t+1)p+q$ with $10\leq q\leq t-$ 1, then
$(p-1)t+p<n-t+1\leq(t+1)p.$
Let $G=\{a_{1},a_{2},\ldots,a_{2p}\}$. Then $n\geq a_{2p}=(t+1)p\geq n+t-1$ and $a_{2p-1}=t(p-1)+p<n-t+1$, and so it is routine to verify that $G\in C_{n,t}$. This proves $D\geq 2p$. To prove $D\leq 2p$, we assume on the contrary that there exists $F=\{i_{1},\ldots,i_{r}\}\in C_{n,t}$ with $1\leq i_{1}<\cdots<i_{r}\leq n$ such that $r>2p$. Then $i_{r-1}\geq i_{2p}\geq a_{2p}\geq n-t+1$. This is contradicted to condition (5). Hence, $D=2p$.
Case 2: If $n=(t+1)p+t$, then $n-t+1=(t+1)p+1=a_{2p+1},$ and one easily checks that $\{a_{1},\ldots,a_{2p+1}\}$ belongs to $C_{n,t}$. It follows that $D=2p+1$, by using the same argument as in case 1.
The theory of Alexander duality plays an important role in the study of squarefree monomial ideals. Let $I=I(\Delta)$ be a squarefree monomial ideal. The Alexander dual of $I$ is defined to be the monomial ideal
$I^{\vee}=(x_{C}\colon\;C\mbox{ is a vertex cover of }\Delta).$
A well-known result states that $S/I$ is Cohen-Macaulay if and only if $I^{\vee}$ has a linear resolution, (see e.g. [10, Theorem 8.1.9]). Moreover, according to [10, Theorem 8.2.20], we know that $S/I$ is sequentially Cohen-Macaulay if and only if $I^{\vee}$ is componentwise linear. Lastly, for any squarefree monomial ideal $I$ of $S$, we have $\operatorname{pd}(S/I)=\operatorname{reg}(I^{\vee})$, as stated in [10, Proposition 8.1.10]. Here, $\operatorname{pd}$ refers to the projective dimension and $\operatorname{reg}$ refers to the (Castelnuovo-Mumford) regularity. To learn more about Alexander duality, one can refer to Chapter 8 of the book [10].
The following result recovers [13, Theorem 4.1], which is also proven in a different manner in [4].
Corollary 2.10.
Let $2\leq t\leq n$ and write $n=(t+1)p+q$ with $0\leq q\leq t$. Then $S/I_{t}(L_{n})$ is sequentially Cohen-Macaulay and
$\mathrm{pd}(S/I_{t}(L_{n}))=\left\{\begin{array}[]{ll}2p,&\hbox{$q<t$;}\\ 2p+1,&\hbox{$q=t$.}\end{array}\right..$
Proof.
In [9, Corollary 5.4], it was shown that if $\Delta$ is a simplicial forest then $S/I(\Delta)$ is sequentially Cohen-Macaulay. Applying this result and in view of Lemma 2.2, we can conclude that $S/I_{t}(L_{n})$ is sequentially Cohen-Macaulay.
By [10, Theorem 8.2.20], we know $I_{t}(L_{n})^{\vee}$ is componentwise linear. Therefore, using [10, Corollary 8.2.14], we have $\operatorname{reg}(I_{t}(L_{n})^{\vee})=\operatorname{Deg}(I_{t}(L_{n})^{\vee})$. Furthermore, according to [10, Proposition 8.1.10], we have $\operatorname{pd}(S/I_{t}(L_{n}))=\operatorname{Deg}(I_{t}(L_{n})^{\vee})$. The result now follows from Corollary 2.9.
3. Multiplicity
In this section, we present the main result of this paper, where we provide an explicit expression for the multiplicity of powers of $I_{t}(L_{n}).$ We begin with a result regarding the multiplicity of square-free monomial ideals.
Lemma 3.1.
Let $I$ be a square-free monomial ideal of height $h$. Then $\operatorname{mult}(S/I)$ is equal to the number of minimal prime ideals of $I$ of height $h$.
Proof.
This follows directly from [10, Corollary 6.2.3].
[…]
Lemma 3.2.
Assume that $n=at+b$ with $a\geq 1$ and $0\leq b<t$. Then
$\operatorname{height}(I_{t}(L_{n}))=a.$
Proof.
[…]
We next describe the prime ideals of $I_{t}(L_{n})$ of height $a$.
Lemma 3.3.
Assume $n=at+b$ with $a\geq 1$ and $0\leq b<t$. Let $\operatorname{\mathfrak{p}}=(x_{i_{1}},\ldots,x_{i_{a}})$ with $1\leq i_{1}<\cdots<i_{a}\leq n$. Then $\operatorname{\mathfrak{p}}\in\mathrm{ass}(I_{t}(L_{n}))$ if and only if the following conditions are satisfied:
  1. (1)
    for $1\leq j\leq a$, one has $(j-1)t+b+1\leq i_{j}\leq jt$;
  2. (2)
    for $11\leq j\leq a-$ 1, one has $i_{j+1}-i_{j}\leq t$.
Proof.
Assume that the elements of $F=\{i_{1},\ldots,i_{a}\}$ satisfies (1), (2). We only need to prove $F\in C_{n,t}$, as defined in Notation 2.6. First, it is easy to see that $i_{1}\leq t$ and $i_{2}\geq(2-1)t+b+1\geq t+1$. Next, we observe that $i_{a-1}\leq(a-1)t=n-t-b<n-t+1$ and $i_{a}\geq(a-1)t+b+1=n-t+1$. Finally, for $j=1,\ldots,a-2$, we see that
$i_{j+2}-i_{j}\geq(j+1)t+b+1-jt=t+b+1\geq t+1.$
Thus, $F\in C_{n,t}$ and so $\operatorname{\mathfrak{p}}\in\mathrm{ass}(I_{t}(L_{n}))$ by Theorem 2.7.
Conversely, first note that $\{i_{1},\ldots,i_{a}\}$ is a minimal vertex cover of $\Delta_{n,t}$. Since the $a$ facets $F_{1},F_{t+1},\ldots,F_{(a-1)t+1}$ of $\Delta_{n,t}$ are pairwise disjoint, it follows that $i_{j}\in F_{j}$ for $j=1,\ldots,a$, and thus $(j-1)t+b+1\leq i_{j}\leq jt$. This implies (1) has to be satisfied. As to condition (2), it follows immediately from Theorem 2.7.
To determine the multiplicity of $S/I_{t}(L_{n})$, in view of Lemma 3.1, we need to count the number of the sequences $(i_{1},\ldots,i_{a})$ that satisfy condition (1) and (2), as stated in Lemma 3.3. This is a purely combinatorial question, which will be solved using the following two results.
Lemma 3.4.
Given $a\geq 1$ and $t\geq 1$, let $T_{a,t}$ denote the set of integral sequences $i_{1}<\ldots<i_{a}$ such that
  1. (1)
    for $1\leq j\leq a$, one has $(j-1)t+1\leq i_{j}\leq jt$;
  2. (2)
    for $11\leq j\leq a-$ 1, one has $i_{j+1}-i_{j}\leq t$.
Then
$|T_{a,t}|=\begin{pmatrix}a+t-1\\ a\end{pmatrix}.$
Proof.
We proceed by induction on $t$. If $t=1$, the result follows easily. Assume now that $t\geq 2$. For each sequence $(i_{1},i_{2},\ldots,i_{a})\in T_{a,t}$, we define
$\tau(i_{1},\ldots,i_{a}):=\operatorname{min}\{j\in[a]\colon\;i_{j}=(j-1)t+1\}.$
By convention, we set $\tau(i_{1},\ldots,i_{a})=a+1$ if $i_{j}\neq(j-1)t+1$ for all $j=1,\ldots,a$. Thus, $\tau(i_{1},\ldots,i_{a})\in\{1,\ldots,a,a+1\}$. We make the following easy observation:
if $\tau(i_{1},\ldots,i_{a})=k<a$, then $i_{j}=(j-1)t+1$ for $j=k,\ldots,a$.
For $k\in[a+1]$, we define
$T_{a,t}^{k}=\{(i_{1},i_{2},\ldots,i_{a})\in T_{a,t}\colon\;\tau(i_{1},\ldots,i_{a})=k\}.$
One easily see that $|T_{a,t}^{1}|=1$ and $|T_{a,t}^{2}|=|\{i_{1}:2\leq i_{1}\leq t\}|=t-1$.
In general, for $k\in[a+1]$, we let $T_{k}$ denote the set of all sequences $i_{1}<i_{2}<\ldots<i_{k-1}$ such that $(j-1)t+2\leq i_{j}\leq jt$ for $j=1,\ldots,k-1,$ and $i_{j+1}-i_{j}\leq t$ for $j=1,\ldots,k-2.$ It follows from the above observation that $|T_{a,t}^{k}|=|T_{k}|$ for $k=1,\ldots,a+1$.
Set $i_{j}^{\prime}=i_{j}-j$ for $j=1,\ldots,k-1$. Then, we have for $j=1,\ldots,k-1$,
$(j-1)t+2\leq i_{j}\leq jt\Longleftrightarrow(j-1)(t-1)+1\leq i_{j}^{\prime}\leq j(t-1);$
and for $j=1,\ldots,k-2$,
$i_{j+1}-i_{j}\leq t\Longleftrightarrow i^{\prime}_{j+1}-i^{\prime}_{j}\leq t-1.$
Thus,
$(i_{1},\ldots,i_{k-1})\in T_{k}\Longleftrightarrow(i_{1}^{\prime},\ldots,i_{k-1}^{\prime})\in T_{t-1,k-1}.$
This implies the following map:
$(i_{1},\ldots,i_{a})\longmapsto(i_{1}^{\prime},\ldots,i_{a}^{\prime})$
establishes a bijection between $T_{k}$ and $T_{t-1,k-1}$. Hence, $|T_{k}|$ is equal to the number of $T_{t-1,k-1}$. By inductive hypothesis, we have $|T_{t-1,k-1}|=\begin{pmatrix}k+t-3\\ k-1\end{pmatrix}.$ Hence, it follows that
$|T_{a,t}|$ $=\mathop{\sum}\limits_{k=1}^{a+1}|T^{k}_{a,t}|=\mathop{\sum}\limits_{k=1}^{a+1}\begin{pmatrix}k+t-3\\ k-1\end{pmatrix}$
$=\begin{pmatrix}t-2\\ 0\end{pmatrix}+\begin{pmatrix}t-1\\ 1\end{pmatrix}+\begin{pmatrix}t\\ 2\end{pmatrix}+\cdots+\begin{pmatrix}a+t-3\\ a\end{pmatrix}+\begin{pmatrix}a+t-2\\ a\end{pmatrix}$
$=(\begin{pmatrix}t-1\\ 0\end{pmatrix}+\begin{pmatrix}t-1\\ 1\end{pmatrix})+\begin{pmatrix}t\\ 2\end{pmatrix}+\cdots+\begin{pmatrix}a+t-3\\ a\end{pmatrix}+\begin{pmatrix}a+t-2\\ a\end{pmatrix}$
$=\begin{pmatrix}t\\ 1\end{pmatrix}+\begin{pmatrix}t\\ 2\end{pmatrix}+\cdots+\begin{pmatrix}a+t-3\\ a\end{pmatrix}+\begin{pmatrix}a+t-2\\ a\end{pmatrix}$
$=\begin{pmatrix}t+a-2\\ a-1\end{pmatrix}+\begin{pmatrix}a+t-2\\ a\end{pmatrix}$
$=\begin{pmatrix}a+t-1\\ a\end{pmatrix}.$
Here, we use the formula $\binom{n}{k}+\binom{n}{k+1}=\binom{n+1}{k+1}$ many times.
Corollary 3.5.
Given $n\geq t\geq 1$ and assume $n=at+b$ with $10\leq b\leq t-$ 1. Let $X_{n,t}$ denote the set of integral sequences $i_{1}<\ldots<i_{a}$ such that
  1. (1)
    for $1\leq j\leq a$, one has $(j-1)t+b+1\leq i_{j}\leq jt$;
  2. (2)
    for $11\leq j\leq a-$ 1, one has $i_{j+1}-i_{j}\leq t$.
Then
$|X_{n,t}|=\begin{pmatrix}a+t-b-1\\ a\end{pmatrix}.$
Proof.
Set $i_{j}^{\prime}=i_{j}-bj$. Then, we have for $1\leq j\leq a$,
$(j-1)t+b+1\leq i_{j}\leq jt\Longleftrightarrow(j-1)(t-b)+1\leq i_{j}^{\prime}\leq j(t-b),$
and for $11\leq j\leq a-$ 1,
$i_{j+1}-i_{j}\leq t\Longleftrightarrow i_{j+1}^{\prime}-i_{j}^{\prime}\leq t-b.$
Thus, the map: $(i_{1},\ldots,i_{a})\longmapsto(i_{1}^{\prime},\ldots,i_{a}^{\prime})$ establishes a bijection between $X_{n,t}$ and $T_{a,t-b}$. This implies
$|X_{n,t}|=|T_{a,t-b}|=\begin{pmatrix}a+t-b-1\\ a\end{pmatrix},$
as desired.
Proposition 3.6.
Assume $n=at+b$ with $a\geq 1$ and $0\leq b<t$. Then
$\operatorname{mult}(S/I_{t}(L_{n}))=\begin{pmatrix}a+t-b-1\\ a\end{pmatrix}.$
Proof.
This result is a combination of Lemma 3.1, Lemma 3.3, and Corollary 3.5.
Based on Proposition 3.6, our main result can be obtained through an induction process. The crucial step in the proof is to identify suitable short exact sequences. In order to do this, we require the following lemmas.
[…]
$(I+J):K=(I:K)+(J:K)=(I:K)+J.$
This fact is also useful in the following proof.
Lemma 3.7.
For $s\geq 2$, we have
  1. (1)
    $I_{t}(L_{n})^{s}:u_{n-t+1}=I_{t}(L_{n})^{s-1}$;
  2. (2)
    Let $(I_{t}(L_{n-1})^{s}:u_{n-t+1},x_{n-t})$ denote the ideal $(I_{t}(L_{n-1})^{s}:u_{n-t+1})+(x_{n-t})$. Then
    $(I_{t}(L_{n-1})^{s}:u_{n-t+1},x_{n-t})=(I_{t}(L_{n-t-1})^{s},x_{n-t}).$
    In particular, if $n\leq 2t$, then $I_{t}(L_{n-t-1})^{s}=0$ and so $(I_{t}(L_{n-1})^{s}:u_{n-t+1},x_{n-t})=(x_{n-t}).$
Proof.
(1) is just [3, Lemma 2.1].
(2) Since $x_{n-t}$ does not divide $u_{n-t+1}$, it follows that
$(I_{t}(L_{n-1})^{s}:u_{n-t+1})+(x_{n-t})=(I_{t}(L_{n-1})^{s},x_{n-t}):u_{n-t+1}.$
It is easy to see that $(I_{t}(L_{n-1})^{s},x_{n-t})=(I_{t}(L_{n-t-1})^{s},x_{n-t}).$ Note that $(I_{t}(L_{n-t-1})^{s},x_{n-t})$ and $u_{n-t+1}$ have disjoint supports, the result follows.
Lemma 3.8.
Let $A_{i}:=(I_{t}(L_{n-i})^{s},x_{n-t+1}\cdots x_{n-i+1})$ and let $B_{i}:=(A_{i},x_{n-i+1})$ for $i=1,\ldots,t$. Then $B_{i}=(I_{t}(L_{n-i})^{s},x_{n-i+1})$ and $A_{i}:x_{n-i+1}=A_{i+1}$ for $i=1,\ldots,t$.
Proof.
The equality that $B_{i}=(I_{t}(L_{n-i})^{s},x_{n-i+1})$ follows easily. Let us show that $A_{i}:x_{n-i+1}=A_{i+1}$. Since
$I_{t}(L_{n-i})^{s}=(I_{t}(L_{n-i-1})^{s},u_{n-i-t+1}I_{t}(L_{n-i})^{s-1}),$
we have
$A_{i}:x_{n-i+1}=(I_{t}(L_{n-i})^{s},x_{n-t+1}\cdots x_{n-i+1}):x_{n-i+1}$
$=(I_{t}(L_{n-i-1})^{s},u_{n-i-t+1}I_{t}(L_{n-i})^{s-1}:x_{n-i+1})+(x_{n-t+1}\cdots x_{n-i})$
$=(I_{t}(L_{n-i-1})^{s},u_{n-i-t+1}I_{t}(L_{n-i})^{s-1},x_{n-t+1}\cdots x_{n-i}):x_{n-i+1}$
$=(I_{t}(L_{n-i-1})^{s},x_{n-t+1}\cdots x_{n-i}):x_{n-i+1}=A_{i+1}.$
Here, the third equality follows because $x_{n-t+1}\cdots x_{n-i}$ divides $u_{n-i-t+1}$.
We are now ready to present the main result of this paper.
Theorem 3.9.
Assume $n=at+b$, with $a\geq 1$ and $0\leq b<t$. Then, for all $s\geq 1$, we have
$\operatorname{mult}(S/I_{t}(L_{n})^{s})=\begin{pmatrix}s+a-1\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b-1\\ a\end{pmatrix}.$
Proof.
We proceed by induction on both $s$ and $n$. The case that $s=1$ follows from Proposition 3.6. If $n=t$, we have
$\operatorname{HS}(S/I_{t}(L_{t})^{s},z)=\frac{1-z^{st}}{(1-z)^{t}}=\frac{1}{(1-z)^{t-1}}\sum_{i=0}^{st-1}z^{i}.$
It follows that $\mathbf{Q}(S/I_{t}(L_{t})^{s},z)=\sum_{i=0}^{st-1}z^{i}$ and so $\operatorname{mult}(S/I_{t}(L_{t})^{s})=\mathbf{Q}(S/I_{t}(L_{t})^{s},1)=st.$ This proves the result holds when $n=t$. Now, assume $s>1$ and $n>t$. To begin, we write $n=at+b$, where $a\geq 1$ and $0\leq b<t$. There are two cases to consider.
Case 1: $b>0$. Consider the following three short exact sequences of graded $S$-modules:
$0\longrightarrow\frac{S}{I_{t}(L_{n})^{s}:u_{n-t+1}}[-t]\longrightarrow\frac{S}{I_{t}(L_{n})^{s}}\longrightarrow\frac{S}{(I_{t}(L_{n})^{s},u_{n-t+1})}\longrightarrow 0,$
$0\longrightarrow\frac{S}{I_{t}(L_{n-1})^{s}:u_{n-t+1}}[-t]\longrightarrow\frac{S}{I_{t}(L_{n-1})^{s}}\longrightarrow\frac{S}{(I_{t}(L_{n-1})^{s},u_{n-t+1})}\longrightarrow 0,$
and
$0\longrightarrow\frac{S}{I_{t}(L_{n-1})^{s}:x_{n-t}u_{n-t+1}}[-1]\longrightarrow\frac{S}{I_{t}(L_{n-1})^{s}:u_{n-t+1}}$
$\longrightarrow\frac{S}{(I_{t}(L_{n-1})^{s}:u_{n-t+1},x_{n-t})}\longrightarrow 0.$
Note that $(I_{t}(L_{n})^{s},u_{n-t+1})=(I_{t}(L_{n-1})^{s},u_{n-t+1}),$ it follows from the additivity of Hilbert series that
(1) $\begin{split}&\operatorname{HS}(\frac{S}{I_{t}(L_{n})^{s}},z)=z^{t}\operatorname{HS}(\frac{S}{I_{t}(L_{n})^{s}:u_{n-t+1}},z)+\operatorname{HS}(\frac{S}{I_{t}(L_{n-1})^{s}},z)\\ &-z^{1+t}\operatorname{HS}(\frac{S}{I_{t}(L_{n-1})^{s}:x_{n-t}u_{n-t+1}},z)\quad-z^{t}\operatorname{HS}(\frac{S}{(I_{t}(L_{n-1})^{s}:u_{n-t+1},x_{n-t})},z).\end{split}$
Since $I_{t}(L_{n})^{s}:u_{n-t+1}=I_{t}(L_{n})^{s-1}$ by Lemma 3.7, we have $\dim\frac{S}{I_{t}(L_{n})^{s}:u_{n-t+1}}=n-a$ and thus
(2) $\operatorname{HS}(\frac{S}{I_{t}(L_{n})^{s}:u_{n-t+1}},z)=\frac{1}{(1-z)^{n-a}}\mathbf{Q}(\frac{S}{I_{t}(L_{n})^{s-1}},z).$
Denote $S_{j}=K[x_{1},\ldots,x_{j}]$ for $j=1,\ldots,n$. Since $n-1=at+(b-1)$ with $b-1\geq 0$, it follows that
$\dim\frac{S_{n-1}}{I_{t}(L_{n-1}^{s})}=n-1-a$
and thus
(3) $\operatorname{HS}(\frac{S}{I_{t}(L_{n-1})^{s}},z)=\frac{1}{1-z}\operatorname{HS}(\frac{S_{n-1}}{I_{t}(L_{n-1})^{s}},z)=\frac{1}{(1-z)^{n-a}}\mathbf{Q}(\frac{S_{n-1}}{I_{t}(L_{n-1}^{s})},z).$
Here, we use the assumption that $b\geq 1$. Note that
$I_{t}(L_{n-1})^{s}:x_{n-t}u_{n-t+1}=I_{t}(L_{n-1})^{s}:u_{n-t}x_{n}=I_{t}(L_{n-1})^{s-1}:x_{n}=I_{t}(L_{n-1})^{s-1},$
we have
(4) $\operatorname{HS}(\frac{S}{I_{t}(L_{n-1})^{s}:x_{n-t}u_{n-t+1}},z)=\frac{1}{(1-z)^{n-a}}\mathbf{Q}(\frac{S_{n-1}}{I_{t}(L_{n-1})^{s-1}}),z).$
If $n>2t$, by Lemma 3.7.(2), we have $(I_{t}(L_{n-1})^{s}:u_{n-t+1},x_{n-t})=(I_{t}(L_{n-t-1})^{s},x_{n-t})$. Since $n-t-1=(a-1)t+b-1$ with $b-1\geq 0$, it follows that
$\dim\frac{S_{n-t-1}}{I_{t}(L_{n-t-1})^{s}}=n-t-1-(a-1)=n-a-t,$
and thus
(5) $\begin{split}\operatorname{HS}(\frac{S}{(I_{t}(L_{n-1})^{s}:u_{n-t+1},x_{n-t})},z)=\frac{1}{(1-z)^{t}}\operatorname{HS}(\frac{S_{n-t-1}}{I_{t}(L_{n-t-1})^{s}},z)\\ =\frac{1}{(1-z)^{n-a}}\mathbf{Q}(\frac{S_{n-t-1}}{I_{t}(L_{n-t-1})^{s}},z)\end{split}$
If $n<2t$, then $a=1$ and so
(6) $\operatorname{HS}(\frac{S}{(I_{t}(L_{n-1})^{s}:u_{n-t+1},x_{n-t})},z)=\operatorname{HS}(\frac{S}{(x_{n-t})},z)=\frac{1}{(1-z)^{n-1}}=\frac{1}{(1-z)^{n-a}}$
If $n>2t$, substitute equations (2), (3), (4), and (5) into equation (1), we obtain
$\begin{split}\mathbf{Q}(\frac{S}{I_{t}(L_{n})^{s}},z)=&z^{t}\mathbf{Q}(\frac{S}{I_{t}(L_{n})^{s-1}},z)+\mathbf{Q}(\frac{S_{n-1}}{I_{t}(L_{n-1})^{s}},z)\\ &-z^{1+t}\mathbf{Q}(\frac{S_{n-1}}{I_{t}(L_{n-1})^{s-1}},z)-z^{t}\mathbf{Q}(\frac{S_{n-t-1}}{I_{t}(L_{n-t-1})^{s}},z).\end{split}$
By substituting $z=1$ into the above equation and using the induction hypothesis, it follows that
$\operatorname{mult}(S/I_{t}(L_{n})^{s})=\operatorname{mult}(S/I_{t}(L_{n})^{s-1})+\operatorname{mult}(S_{n-1}/I_{t}(L_{n-1})^{s})-\operatorname{mult}(S_{n-1}/I_{t}(L_{n-1})^{s-1})$
$\qquad\qquad-\operatorname{mult}(S_{n-t-1}/I_{t}(L_{n-t-1})^{s})$
$=\begin{pmatrix}s+a-2\\ s-2\end{pmatrix}\begin{pmatrix}a+t-b-1\\ a\end{pmatrix}+\begin{pmatrix}s+a-1\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b\\ a\end{pmatrix}$
$~{}~{}~{}~{}-\begin{pmatrix}s+a-2\\ s-2\end{pmatrix}\begin{pmatrix}a+t-b\\ a\end{pmatrix}-\begin{pmatrix}s+a-2\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b-1\\ a-1\end{pmatrix}$
$=\begin{pmatrix}s+a-1\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b\\ a\end{pmatrix}-\begin{pmatrix}s+a-2\\ s-2\end{pmatrix}\begin{pmatrix}a+t-b-1\\ a-1\end{pmatrix}-\begin{pmatrix}s+a-2\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b-1\\ a-1\end{pmatrix}$
$=\begin{pmatrix}s+a-1\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b\\ a\end{pmatrix}-\begin{pmatrix}s+a-1\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b-1\\ a-1\end{pmatrix}$
$=\begin{pmatrix}s+a-1\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b-1\\ a\end{pmatrix}.$
If $n<2t$, by substituting equations (2), (3), (4), and (6) into equation (1), we obtain
$\begin{split}\mathbf{Q}(\frac{S}{I_{t}(L_{n})^{s}},z)=&z^{t}\mathbf{Q}(\frac{S}{I_{t}(L_{n})^{s-1}},z)+\mathbf{Q}(\frac{S_{n-1}}{I_{t}(L_{n-1})^{s}},z)\\ &-z^{1+t}\mathbf{Q}(\frac{S_{n-1}}{I_{t}(L_{n-1})^{s-1}},z)-z^{t}\end{split}$
By substituting $z=1$ into the above equation and using the induction hypothesis, it follows that:
$\operatorname{mult}(S/I_{t}(L_{n})^{s})$
$=\operatorname{mult}(S/I_{t}(L_{n})^{s-1})+\operatorname{mult}(S_{n-1}/I_{t}(L_{n-1})^{s})-\operatorname{mult}(S_{n-1}/I_{t}(L_{n-1})^{s-1})-1$
$=\begin{pmatrix}s-1\\ s-2\end{pmatrix}\begin{pmatrix}t-b\\ 1\end{pmatrix}+\begin{pmatrix}s\\ s-1\end{pmatrix}\begin{pmatrix}t-b+1\\ 1\end{pmatrix}-\begin{pmatrix}s-1\\ s-2\end{pmatrix}\begin{pmatrix}t-b+1\\ 1\end{pmatrix}-1$
$=\begin{pmatrix}s\\ s-1\end{pmatrix}\begin{pmatrix}t-b+1\\ 1\end{pmatrix}-\begin{pmatrix}s-1\\ s-2\end{pmatrix}\begin{pmatrix}t-b\\ 0\end{pmatrix}-\begin{pmatrix}s-1\\ s-1\end{pmatrix}\begin{pmatrix}t-b\\ 0\end{pmatrix}$
$=\begin{pmatrix}s\\ s-1\end{pmatrix}\begin{pmatrix}t-b+1\\ 1\end{pmatrix}-\begin{pmatrix}s\\ s-1\end{pmatrix}\begin{pmatrix}t-b\\ 0\end{pmatrix}$
$=\begin{pmatrix}s\\ s-1\end{pmatrix}\begin{pmatrix}t-b\\ 1\end{pmatrix}.$
Case 2: $b=0$. We may assume $a>1$, since the case that $a=1$, i.e, $n=t$ has been addressed. Put $A_{i}:=(I_{t}(L_{n-i})^{s},x_{n-t+1}\cdots x_{n-i+1})$ and $B_{i}:=(A_{i},x_{n-i+1})$ for $i=1,\ldots,t$. According to Lemma 3.8, we obtain the following $t-1$ short exact sequences:
$0\longrightarrow\frac{S}{A_{2}}[-1]\longrightarrow\frac{S}{A_{1}}\longrightarrow\frac{S}{B_{1}}\longrightarrow 0,$
$0\longrightarrow\frac{S}{A_{3}}[-1]\longrightarrow\frac{S}{A_{2}}\longrightarrow\frac{S}{B_{2}}\longrightarrow 0,$
$\vdots$
$0\longrightarrow\frac{S}{A_{t}}[-1]\longrightarrow\frac{S}{A_{t-1}}\longrightarrow\frac{S}{B_{t-1}}\longrightarrow 0$
Since $A_{1}=(I_{t}(L_{n-1})^{s},u_{n-t+1})$, we can use the first short exact sequence used in the proof of the first case and the short exact sequences above to obtain
$\operatorname{HS}(S/I_{t}(L_{n})^{s},z)=z^{t}\operatorname{HS}(S/I_{t}(L_{n})^{s-1},z)+\operatorname{HS}(S/A_{1},z)$
$=z^{t}\operatorname{HS}(S/I_{t}(L_{n})^{s-1},z)+z\operatorname{HS}(S/A_{2},z)+\operatorname{HS}(S/B_{1},z)$
$=z^{t}\operatorname{HS}(S/I_{t}(L_{n})^{s-1},z)+z^{2}\operatorname{HS}(S/A_{3},z)+\operatorname{HS}(S/B_{1},z)+z\operatorname{HS}(S/B_{2},z)$
=$\cdots$= ⋯
$=z^{t}\operatorname{HS}(S/I_{t}(L_{n})^{s-1},z)+z^{t-1}\operatorname{HS}(S/A_{t},z)+\mathop{\sum}\limits_{i=1}^{t-1}z^{i-1}\operatorname{HS}(S/B_{i},z).$
Since $A_{t}=B_{t}$ by their definitions, it follows that
(7) $\operatorname{HS}(S/I_{t}(L_{n})^{s},z)=z^{t}\operatorname{HS}(S/I_{t}(L_{n})^{s-1},z)+\mathop{\sum}\limits_{i=1}^{t}z^{i-1}\operatorname{HS}(S/B_{i},z)$
Take any $i\in\{1,\ldots,t\}$. By Lemma 3.8, we see that $B_{i}=(I_{t}(L_{n-i})^{s},x_{n-i+1})$. Since $n-i=(a-1)t+t-i$, it follows $\dim S_{n-i}/I_{t}(L_{n-i})^{s}=n-i-a+1$, and so
(8) $\operatorname{HS}(S/B_{i},z)=\frac{1}{(1-z)^{i-1}}\operatorname{HS}(S_{n-i}/I_{t}(L_{n-i})^{s},z)=\frac{\mathbf{Q}(S_{n-i}/I_{t}(L_{n-i})^{s},z)}{(1-z)^{n-a}}$
Substitute equation (8) into equation (7), we have
(9) $\mathbf{Q}((S/I_{t}(L_{n})^{s},z)=z^{t}\mathbf{Q}((S/I_{t}(L_{n})^{s-1},z)+\mathop{\sum}\limits_{i=1}^{t}z^{i-1}\mathbf{Q}(S_{n-i}/I_{t}(L_{n-i})^{s},z)$
This implies
$\operatorname{mult}(S/I_{t}(L_{n})^{s})$
$=\operatorname{mult}(S/I_{t}(L_{n})^{s-1})+\mathop{\sum}\limits_{i=1}^{t}\operatorname{mult}(S_{n-i}/I_{t}(L_{n-i})^{s})$
$=\begin{pmatrix}s+a-2\\ s-2\end{pmatrix}\begin{pmatrix}a+t-1\\ a\end{pmatrix}+\begin{pmatrix}s+a-2\\ s-1\end{pmatrix}\mathop{\sum}\limits_{i=1}^{t}\begin{pmatrix}a+i-2\\ a-1\end{pmatrix}$
$=\begin{pmatrix}s+a-2\\ s-2\end{pmatrix}\begin{pmatrix}a+t-1\\ a\end{pmatrix}+\begin{pmatrix}s+a-2\\ s-1\end{pmatrix}\begin{pmatrix}a+t-1\\ a\end{pmatrix}$
$=\begin{pmatrix}s+a-1\\ s-1\end{pmatrix}\begin{pmatrix}a+t-1\\ a\end{pmatrix}.$
Remark 3.10.
According to Theorem 3.9, it follows that the multiplicity of $S/I_{t}(L_{n})^{s}$ is a polynomial of degree $a$ in $s$.
Acknowledgment: This research is supported by NSFC (No. 11971338)